15-th Conference

of the

International Linear Algebra Society

Westin Resort and Spa Cancún, México June 16–20, 2008

Organizing Committee

Rafael Bru Luz María de Alba Daniel Hershkowitz

Andre Klein Beatrice Meini Dale Olesky

Vadim Olshevsky Jeff Stuart Daniel Szyld

Luis Verde-Star

Local Organizing Committee

María José Arroyo, Universidad Autónoma Metropolitana, Iztapalapa

Rubén Martínez-Avendaño, Universidad Autónoma del Estado de Hidalgo

Martha Takane, Instituto de Matemáticas, Universidad Nacional Autónoma de México

Luis Verde-Star (Chair), Universidad Autónoma Metropolitana, Iztapalapa

Acknowledgements

The organizers gratefully acknowledge the support received from:

Consejo Nacional de Ciencia y Tecnología Rectoría General de la Universidad Autónoma Metropolitana Rectoría de la UAM-Iztapalapa División de Ciencias Básicas e Ingeniería, UAM-Iztapalapa Departamento de Matemáticas, UAM-Iztapalapa Instituto de Matemáticas, Universidad Nacional Autónoma de México Dirección General de Asuntos del Personal Académico, UNAM (Proyectos PAPIIT y PAEP) Universidad Autónoma del Estado de Hidalgo International Linear Algebra Society Taylor and Francis

Personal Schedule

Monday 8:10 Opening ceremony

8:20 L. Hogben 9:15 Mini-break 9:20 E. Jiang

10:15 Coffee break 10:45 11:10 11:35 12:00 12:25

12:50–15:30 Lunch break

15:30 A. Boettcher 16:25 Coffee break

16:55 17:20 17:45 18:10 18:35

Tuesday 8:10 N. Shaked-Monderer 9:05 Mini-break 9:10 D. Kressner 10:05 Coffee break 10:35 11:00 11:25 11:50 12:15 12:40–15:30 Lunch break 15:30 F. M. Dopico 16:25 Coffee break 16:55 17:20 17:45 18:10 18:35

Wednesday 8:10 J. M. Peña 9:05 Mini-break 9:10 J. Nagy 10:05 Coffee break 10:35 11:00 11:25 11:50 12:15 12:40–13:30 Lunch break 13:30–19:00 Excursion to Tulum

Thursday 8:10 H. Fassbender 9:05 Mini-break 9:10 P. Van Dooren 10:05 Coffee break 10:35 11:00 11:25 11:50 12:15 12:40–14:00 Lunch break 14:00 ILAS Business Meeting 15:30 L. Gemignani 16:25 Coffee break 16:55 17:20 17:45 18:10 18:35

20:00 Banquet

Friday 8:10 P. Rosenthal 9:05 Mini-break 9:10 I. Spitkovsy 10:05 Coffee break 10:35 11:00 11:25 11:50 12:15 12:40–15:30 Lunch break 15:30 15:55 16:20 16:45 17:10 17:35 Closing Ceremony

Schedule for ILAS, Cancún, June 16–20, 2008

Morning sessions

Monday 8:10–8:20

Opening ceremony 8:20–9:15 IPL

L. Hogben

9:15– 9:20 Mini-break

9:20–10:15 IPL

E. Jiang 10:15–10:45 Coffee break 10:45– 12:50

Parallel session 1 5 talks of 25min MS4, MS7, CT 12:50–15:30

Lunch break

Tuesday 8:10–9:05 IPL

N. Shaked-Monderer

9:05–9:10 Mini-break 9:10–10:05 IPL

D. Kressner

10:05–10:35 Coffee break 10:35–12:40

Parallel session 3 5 talks of 25min MS6, MS8, CT 12:40–15:30

Lunch break Wednesday 8:10–9:05 IPL

J. M. Peña

9:05–9:10 Mini-break 9:10–10:05 IPL

J. Nagy

10:05–10:35 Coffee break 10:35–12:40

Parallel session 5 5 talks of 25min

MS1, MS4, MS7, CT 12:40–13:30

Lunch break 13:30–19:00

Excursion to Tulum

Thursday 8:10–9:05 IPL

H. Fassbender

9:05–9:10 Mini-break 9:10–10:05 IPL

P. Van Dooren

10:05–10:35 Coffee break 10:35–12:40

Parallel session 6 5 talks of 25min MS1, MS5, CT 12:40–14:00

Lunch break

Friday 8:10–9:05 IPL

P. Rosenthal

9:05–9:10 Mini-break 9:10–10:05 IPL

I. Spitkovsy

10:05–10:35 Coffee break 10:35–12:40

Parallel session 8 5 talks of 25min MS2, MS3, CT 12:40–15:30

Lunch break

Afternoon sessions

Monday

15:30–16:25 IPL

A. Boettcher

16:25–16:55 Coffee break 16:55–19:00

Parallel session 2 5 talks of 25min

MS8, CT

Tuesday

15:30–16:25 IPL

F. M. Dopico

16:25–16:55 Coffee break 16:55–19:00

Parallel session 4 5 talks of 25min MS4, MS6, MS7,CT Wednesday

Excursion

to Tulum Thursday 14:00–15:25 ILAS Business Meeting 15:30–16:25 IPL

L. Gemignani

16:25–16:55 Coffee break 16:55–19:00

Parallel session 7 5 talks of 25min MS2, MS3, CT

20:00 ILAS Banquet

Friday

15:30–17:35

Parallel session 9 5 talks of 25min MS1, MS5, CT

17:35–17:45

Closing ceremony

IPL means Invited Plenary Talk. (Salón Pintores)

MS1 Combinatorial Matrix Theory; Parallel Sessions 5, 6, 9. MS2 Eigenproblems: Theory and computation; Parallel Sessions 7, 8. MS3 Implementation and Application issues in regularizing least squares and total least squares; Parallel Sessions 7, 8. MS4 Linear Algebra Education; Parallel Sessions 1, 4, 5. MS5 Linear Algebra in Model Reduction; Parallel Sessions 6, 9. MS6 Matrix functions and matrix equations; Parallel Sessions 3, 4. MS7 Max Algebra; Parallel Sessions 1, 4, 5. MS8 Nonnegative and eventually nonnegative matrices; Parallel Sessions 2, 3. CT Contributed talks, all Parallel Sessions

Invited plenary lectures

• Monday

– 8:20–9:15 Leslie Hogben (Ames, Iowa, USA)

Minimum Rank Problems: Recent Developments

– 9:20–10:15 ErXiong Jiang (Shanghai, China)

Some inverse eigenvalue problems for Jacobi matrices

– 15:30–16:25 Albrecht Boettcher (Chemnitz, Germany)

Toeplitz matrices with Fisher-Hartwig symbols

• Tuesday

– 8:10–9:05 Naomi Shaked-Monderer (Emek Yezreel, Israel) Completely Positive Matrices and the CP-rank

– 9:10–10:05 Daniel Kressner (Zurich, Switzerland)

Matrix product eigenvalue problems

– 15:30–16:25 Froilán M. Dopico (Madrid, Spain)

Implicit Jacobi algorithms for the symmetric eigenproblem

• Wednesday

– 8:10–9:05 Juan Manuel Peña (the LAMA Lecture) (Zaragoza, Spain) From Total Positivity to Positivity: related classes of matrices

– 9:10–10:05 James Nagy (the SIAM Lecture) (Atlanta, Georgia, USA) Kronecker Products in Imaging Sciences

• Thursday

– 8:10–9:05 Heike Fassbender (Braunschweig, Germany)

TBA

– 9:10–10:05 Paul Van Dooren (Louvain, Belgium)

Some graph optimization problems in data mining

– 15:30–16:25 Luca Gemignani (Pisa, Italy)

Eigenvalue Problems for Rank-structured Matrices

• Friday

– 8:10–9:05 Peter Rosenthal (Toronto, Canada)

Invariant subspaces of semigroups of matrices

– 9:10–10:05 Ilya Spitkovsky (Williamsburg, Virginia, USA)

On the current state of the factorization problem for almost periodic matrix functions

Mini Symposia

MS1 Combinatorial Matrix Theory

Organizers: Leslie Hogben, Iowa State University, and Bryan Shader, University of Wyoming. Parallel Sessions 5, 6, 9.

MS2 Eigenproblems: Theory and computation

Organizer: Julio Moro, Universidad Carlos III de Madrid Parallel Sessions 7, 8.

MS3 Implementation and Application issues in regularizing least squares and total least squares

Organizer: Rosie Renaut, Arizona State University Parallel Sessions 7, 8.

MS4 Linear Algebra Education

Organizers: María Trigueros, Instituto Tecnológico Autónomo de México, and Asuman Oktaç, Centro de Investigación y Estudios Avanzados, IPN, México, and PUCV, Chile.

Parallel Sessions 1, 4, 5.

MS5 Linear Algebra in Model Reduction

Organizer: Peter Benner, Chemnitz University of Technology, Germany Parallel Sessions 6, 9.

MS6 Matrix functions and matrix equations

Organizers: Peter Lancaster, University of Calgary, Canada, and Chun-Hua Guo, University of Regina, Canada

Parallel Sessions 3, 4.

MS7 Max Algebra

Organizers: Hans Schneider, University of Wisconsin, and Peter Butkovic, University of Birmingham, UK Parallel Sessions 1, 4, 5.

MS8 Nonnegative and eventually nonnegative matrices

Organizer: Judi McDonald, Washington State University, USA Parallel Sessions 2, 3.

Parallel Session 1, Monday Morning

Time  MS4 Room 1  MS7 Room 2  CT Room 3  CT Room 4
10:45  Harel, Guershon  Gaubert, Stephane  Dodig, Marija  Costa, Liliana
11:10 Dogan-Dunlap, Hamide van den Driessche, Pauline Corral, Cristina Fernandes, Rosário
11:35  Maracci, Mirko  McEneaney, William  Fošner, Ajda  Stuart, Jeffrey
12:00  Parraguez, Marcela  Zimmermann, Karel  Horn, Roger  Weaver, James
12:25  Trigueros, María  Ponce, Daniela  Stosic, Marko  Vallejo, Ernesto

MS4 Linear Algebra Education, Monday Morning, Room 1

• 10:45 Intellectual Need and Its Role in Mathematics Instruction: Focus on Linear Algebra, Harel, Guershon, University of California, San Diego, USA

• 11:10 Thinking Modes Revealed on Students’ Responses from an Assignment on Linear Independence, Dogan-Dunlap, Hamide, University of Texas at El Paso, USA

• 11:35 Basic notions of Vector Space Theory: students’ models and conceptions, Maracci, Mirko, Siena University, Italy

• 12:00 Construction of a vector space schema, Parraguez, Marcela, PUCV, Chile, (with Oktaç, Asuman)

• 12:25 Spanning sets and vector spaces they generate: an APOS analysis, Trigueros, María, Instituto Tecnológico Autónomo de México, (with Ku, Darly and Oktaç, Asuman)

MS7 Max Algebra, Monday Morning, Room 2

• 10:45 Using max-plus eigenvalues to bound the roots of a polynomial, Gaubert, Stephane, INRIA, Paris, France, (with Marianne Akian (INRIA)and Adrien Brandejsky (ENS Cachan))

• 11:10 Bounds for the Perron root using max eigenvalues, van den Driessche, Pauline, University of Victoria, Canada, (with Elsner, Ludwig)

• 11:35 Max-Plus Bases, Cornices and Pruning, McEneaney, William, University of California, San Diego, USA • 12:00 Solving two-sided (max,plus)-linear equation systems,

Zimmermann, Karel, Charles University, Prague, Czech Republic

• 12:25 NP-hard problems in extremal algebras tackled by particle swarm optimization, Ponce, Daniela, University of Hradec Kralove, Czech Republic, (with Gavalec, Martin)

Contributed Talks, Monday Morning, Room 3

• 10:45 Singular systems, state feedback problem Dodig, Marija, CELC, Universidade de Lisboa, Lisbon, Portugal

• 11:10 On Schur complements of H-matrices Corral, Cristina, Universidad Politécnica de Valencia, Valencia, Spain (with R. Bru, I. Giménez, J. Mas)

• 11:35 Commutativity preserving maps on real matrices Fošner, Ajda, Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

• 12:00 A Canonical Form for Quasi-Real Normal Matrices Under Real Orthogonal Similarity Horn, Roger, University of Utah, Salt Lake City, USA (with Geoffrey R. Goodson)

• 12:25 On Generalized Procrustes Problem Stosic, Marko, Instituto de Sistemas e Robotica, IST, Lisbon, Portugal (with João Xavier)

Contributed Talks, Monday Morning, Room 4

• 10:45 Acyclic Birkhoff Polytope Costa, Liliana, Centre for Research on Optimization and Control, Aveiro, Portugal (with C. M. da Fonseca and Enide Andrade Martins)

• 11:10 Rank partitions and covering numbers under small perturbations of an element Fernandes, Rosário, Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal

• 11:35 Spectrally Arbitrary Ray Patterns Stuart, Jeffrey, Pacific Lutheran University, Tacoma, USA

• 12:00 Nonsingularity of Divisor Tournaments Weaver, James, University of West Florida, Pensacola, FL 32514, USA (with Rohan Hemasinha and Jeffrey L. Stuart)

• 12:25 Additivity obstructions for integral matrices and pyramids Vallejo, Ernesto, Instituto de Matemáticas, Morelia, México (with Miguel Santoyo)

Parallel Session 2, Monday Afternoon

Time  MS8 Room 1  CT Room 2  CT Room 3  CT Room 4
16:55 McDonald, Judith  Machado, Silvia  Gouveia, María  Furuichi, Shigeru
17:20 Šmigoc, Helena  Poole, George  Domínguez, María Elena Bourgeois, Gerald
17:45 Laffey, Thomas  Vargas, Xaab Nop  Rump, Siegfried M.  Soares, Graça
18:10 Eubanks, Sherod Kopparty, Bhaskara Rao  Bru, Rafael  Uchiyama, Mitsuru
18:35 Catral, Minerva  Vieira, Luis  Grudsky, Sergey  Lee, Hosoo

MS8 Nonnegative and eventually nonnegative matrices, Monday Afternoon, Room 1

• 16:55 Nonnegative and Eventually Nonnegative Matrices, McDonald, Judith, Washington State University, USA

• 17:20 An example of constructing a nonnegative matrix with given spectrum, Šmigoc, Helena, University College Dublin, (with Laffey, Thomas)

• 17:45 Some constructive techniques in the nonnegative inverse eigenvalue problem, Laffey, Thomas, University College Dublin, (with Šmigoc, Helena)

• 18:10 Generalized Soules Matrices, Eubanks, Sherod, Washington State University, USA

• 18:35 The Kemeny Constant in Finite Homogeneous Ergodic Markov Chains, Catral, Minerva, University of Victoria, Canada

Contributed Talks, Monday Afternoon, Room 2

• 16:55 GPEA’s researches about the meta resources in teaching and learning the notion of basis of a vector space Machado, Silvia, Pontificia Universidade Catolica de São Paulo, São Paulo, Brasil (with Bianchini, B. L. and Maranhão, M. C. S. A.)

• 17:20 Whatever Happened to Rook’s Pivoting?

Poole, George, East Tennessee State University, Johnson City, USA

• 17:45 Students difficulties with concept of vector space from point of view of APOS Theory Vargas, Xaab Nop, ICYTDF, México, México • 18:10 Generalized inverses of infinite matrices

Kopparty, Bhaskara Rao, Indiana State University, Terre Haute, IN, USA

• 18:35 Euclidean Jordan algebras and inequalities on the parameters and on the spectra of a strongly regular graph Vieira, Luis, Feup, Porto, Portugal

Contributed Talks, Monday Afternoon, Room 3

• 16:55 On a singular Toeplitz pencil Gouveia, María, Department of Mathematics, FCTUC, Coimbra, Portugal

• 17:20 General solution of certain matrix equations arising in filter design applications Domínguez, María Elena, Universidad Politécnica de Madrid, Madrid, Spain

• 17:45 The ratio between the Toeplitz and the unstructured condition number Rump, Siegfried M., Hamburg University of Technology, Hamburg, Germany (with H. Sekigawa)

• 18:10 On some classes of H-matrices Bru, Rafael, Univ. Politécnica, Valencia, Spain (with Ljiljana Cvetković, Vladimir Kostić and Francisco Pedroche)

• 18:35 Uniform boundedness of Toeplitz Matrices with variable coefficients Grudsky, Sergey, CINVESTAV del I.P.N., México, México

Contributed Talks, Monday Afternoon, Room 4

• 16:55 On trace inequalities for products of matrices Furuichi, Shigeru, Nihon University, Tokyo, Japan

• 17:20 About the logarithm function over the matrices Bourgeois, Gerald, Faculté de Luminy, Marseille, France

• 17:45 Inequalities on an indefinite inner product space Soares, Graça, University of Tras-os-Montes and Alto Douro, Vila Real, Portugal (with N. Bebiano et al.)

• 18:10 A New Majorization between functions Uchiyama, Mitsuru, Shimane University, Matsue, Shimane, Japan

• 18:35 Contractions and nonlinear matrix equations on positive definite cones Lee, Hosoo, Kyungpook National University, Daegu, Korea (with Yongdo Lim)

Parallel Session 3, Tuesday Morning

Time MS6 Room 1  MS8 Room 2  CT Room 3  CT Room 4
10:35  Bini, Dario  Szyld, Daniel  Arav, Marina  Cravo, Glória
11:00 Fassbender, Heike Morris, DeAnne Pruneda, Rosa E.  Peña, Marta
11:25 Guo, Chun-Hua  Szyld, Daniel  Semrl, Peter  Roca, Alicia
11:50 Iannazzo, Bruno Neumann, Michael  Glebsky, Lev  Carriegos, Miguel
12:15  Klein, Andre  Gassó, Maria T.  Ferrer, Josep

MS6 Matrix functions and matrix equations, Tuesday Morning, Room 1

• 10:35 Fast solution of a certain Riccati Equation through Cauchy-like matrices, Bini, Dario, University of Pisa, Italy, (with Beatrice Meini and Federico Poloni)

• 11:00 On the numerical solution of large-scale sparse discrete-time Riccati equations, Fassbender, Heike, TU Braunschweig, Germany, (with Peter Benner)

• 11:25 On Newton’s method and Halley’s method for p-th roots of matrices, Guo, Chun-Hua, University of Regina, Canada

• 11:50 Matrix iterations for matrix functions, Iannazzo, Bruno, Università dell’Insubria, Como, Italy

• 12:15 Tensor Sylvester matrices and information matrices of multiple stationary processes , Klein, Andre, University of Amsterdam, The Netherlands

MS8 Nonnegative and eventually nonnegative matrices, Tuesday Morning, Room 2

• 10:35 On General Matrices Having the Perron-Frobenius Property, Szyld, Daniel, Temple University, USA, (with Elhashash, Abed)

• 11:00 Jordan forms corresponding to nonnegative and eventually nonnegative matrices, Morris, DeAnne, Washington State University, USA, (with McDonald, Judith)

• 11:25 Convergence of Stationary Iterative Methods for Hermitian Semidefinite Linear Systems, Szyld, Daniel, Temple University, USA, (with Frommer, Andreas and Nabben, Reinhard)

• 11:50 On Optimal Condition Numbers For Markov Chains, Neumann, Michael, University of Connecticut, USA, (with Sze, Nung-Sing)

Contributed Talks, Tuesday Morning, Room 3

• 10:35 Sign Patterns That Require Almost Unique Rank Arav, Marina, Georgia State University, Atlanta, GA, USA (with Frank Hall, Zhongshan Li, Assefa Merid, Yubin Gao)

• 11:00 Complete Orthogonal Decomposition Compared with Direct Projection Methods Pruneda, Rosa E., University of Castilla-La Mancha, Ciudad Real, Spain (with Beatriz Lacruz)

• 11:25 Locally linearly dependent operators Semrl, Peter, University of Ljubljana, Ljubljana, Slovenia

• 11:50 On low rank perturbations of matrices Glebsky, Lev, Universidad Autónoma de San Luis Potosi, San Luis Potosi, México (with Luis Manuel Rivera)

• 12:15 The class of Inverse-Positive matrices with checkerboard pattern Gassó, Maria T., Inst. Mat. Mult. Universidad Politécnica Valencia, Valencia, Spain (with Manuel F. Abad, and Juan R. Torregrosa)

Contributed Talks, Tuesday Morning, Room 4

• 10:35 Controllability of Matrices with Prescribed Blocks Cravo, Glória, University of Madeira and CELC, Funchal, Portugal

• 11:00 Perturbations preserving conditioned invariant subspaces Peña, Marta, Universitat Politecnica de Catalunya, Barcelona, Spain (with A. Compta and J. Ferrer)

• 11:25 Pencils with Prescribed Constant Subpencils Roca, Alicia, Dpto. de Matemática Aplicada, Universidad Politécnica, Valencia, Spain (with F. C. Silva)

• 11:50 Reachability of regular switched linear systems Carriegos, Miguel, Universidad de León, León, Spain

• 12:15 Geometric structure of the equivalence classes of a controllable pair Ferrer, Josep, Universitat Politecnica de Catalunya, Barcelona, Spain (with A. Compta and M. Peña)

Parallel Session 4, Tuesday Afternoon

Time  MS4 Room 1  MS6 Room 2  MS7 Room 3  CT Room 4
16:55 Klasa-Bompoint, Jacqueline  Lancaster, Peter  Boimond, Jean-Louis  Russo, Maria Rosaria
17:20  Leon, Steven  Meini, Beatrice  Sergeev, Sergey  Bueno Cachadina, M. Isabel
17:45  Tanguay, Denis  Mena, Hermann  Merlet, Glenn  Frank, Martin
18:10  Martin, William  Zúñiga, Juan Carlos  Butkovic, Peter  Salam, Ahmed
18:35  Group discussion  Martínez, José-Javier

MS4 Linear Algebra Education, Tuesday Afternoon, Room 1

• 16:55 Few pedagogical scenarios in Linear Algebra with Cabri and Maple, Klasa-Bompoint, Jacqueline, Dawson College, Canada

• 17:20 When MATLAB Gives ”Wrong” Answers, Leon, Steven, University of Massachusetts Dartmouth, USA

• 17:45 A fundamental paradox in learning algebra, Tanguay, Denis, UQAM, Montreal, Canada, (with Corriveau, Claudia)

• 18:10 Integrating learning theories and application-based modules in teaching linear algebra, Martin, William, North Dakota State University, (with S. Loch, L. Cooley, S. Dexter and D. Vidakovic)

• 18:35 Group discussion

MS6 Matrix functions and matrix equations, Tuesday Afternoon, Room 2

• 16:55 Linearization of Matrix Polynomials, Lancaster, Peter, University of Calgary, Canada

• 17:20 From algebraic Riccati equations to unilateral quadratic matrix equations: old and new algorithms, Meini, Beatrice, University of Pisa, Italy, (with Bini, Dario and Poloni, Federico)

• 17:45 Exponential Integrators for Solving Large-Scale Differential Riccati Equations, Mena, Hermann, Escuela Politécnica Nacional, Quito, Ecuador, (with Benner, Peter)

• 18:10 Matrix polynomials, rational matrices and linear systems: A review, Zúñiga, Juan Carlos, University of Calgary, Canada, (with Lancaster, Peter)

MS7 Max Algebra, Tuesday Afternoon, Room 3

• 16:55 On Steady State Controller in Min-Plus Algebra, Boimond, Jean-Louis, University Angers, France, (with S. Lahaye)

• 17:20 On Kleene stars and intersection of finitely generated semimodules, Sergeev, Sergey, University of Birmingham, UK

• 17:45 Semi-group of matrices acting on the max-plus projective space, Merlet, Glenn, LIAFA CNRS-Universite Paris-Diderot, France

• 18:10 On the permuted max-algebraic eigenvector problem, Butkovic, Peter, University of Birmingham, UK

Contributed Talks, Tuesday Afternoon, Room 4

• 16:55 On some general determinantal identities of Sylvester type Russo, Maria Rosaria, Department of Mathematics - University of Padua, Padova, Italy (with Michela Redivo-Zaglia)

• 17:20 Algorithms for computing the Geronimus Transformation Bueno Cachadina, María Isabel, The University of California at Santa Barbara, Santa Barbara, USA

• 17:45 An iterative method for transport equations in radiotherapy Frank, Martin, University of Kaiserslautern, Kaiserslautern, Germany (with Bruno Dubroca)

• 18:10 A structure-preserving Arnoldi-like method for a class of structured matrices Salam, Ahmed, Université du Littoral-Côte d’Opale, Calais, France

• 18:35 Polynomial regression in the Bernstein basis Martínez, José-Javier, Universidad de Alcalá, Alcalá de Henares, Spain (with Ana Marco)

Parallel Session 5, Wednesday Morning

Time  MS1 Room 1  MS4 Room 2  MS7 Room 3  CT Room 4
10:35 Brualdi, Richard A.  Moura, Ana  Singer, Ivan  Uhlig, Frank
11:00 Hershkowitz, Danny Zandieh, Michelle Loiseau, Jean Jacques Schaffrin, Burkhard
11:25  Kirkland, Steve  Possani, Edgar  Gavalec, Martin  Wu, Pei Yuan
11:50  Pryporova, Olga  Strong, David  Plavka, Jan  Protasov, Vladimir
12:15  Berman, Avi  Group discussion  Schneider, Hans  Seddighin, Morteza

MS1 Combinatorial Matrix Theory, Wednesday Morning, Room 1

• 10:35 A Conjecture in Combinatorial Matrix Theory, Brualdi, Richard A., University of Wisconsin, Madison, USA

• 11:00 On nonnegative sign equivalent and sign similar factorizations of matrices Hershkowitz, Danny, Technion, Israel (with Allan Pinkus)

• 11:25 Constructing Laplacian Integral Split Graphs, Kirkland, Steve, University of Regina, Canada, (with N. Abreu, M. de Freitas and R. Del Vecchio)

• 11:50 Potential Diagonal and D-convergence, Pryporova, Olga, Iowa State University, Ames, IA, USA

• 12:15 The Colin de Verdière Parameter- a progress report, Berman, Avi, Technion, Tel Aviv, Israel

MS4 Linear Algebra Education, Wednesday Morning, Room 2

• 10:35 Skills, Concepts and Models in a Linear Algebra Course, Moura, Ana, Instituto Superior Técnico, UTL, Lisbon, Portugal, (with Santos, P. A.)

• 11:00 Design of a unit to teach eigenvectors and eigenvalues based on the instructional design principles of Realistic Mathematics Education, Zandieh, Michelle, Arizona State University, (with Rasmussen, C. and Larson, C.)

• 11:25 Use of models in the teaching of linear algebra, Possani, Edgar, Instituto Tecnológico Autónomo de México, (with Preciado, G. and Lozano, D.)

• 11:50 A Java Applet and Introductory Tutorial for the Jacobi, Gauss-Seidel and SOR Methods, Strong, David, Pepperdine University, CA, USA

• 12:15 Group discussion

MS7 Max Algebra, Wednesday Morning, Room 3

• 10:35 Max-min convexity, Singer, Ivan, Institute of Mathematics, Bucharest, Romania

• 11:00 Robust stability of positive difference equations, Loiseau, Jean Jacques, University Nantes, France, (with M. Di Loreto)

• 11:25 Permuted max-min eigenvector problem, Gavalec, Martin, University of Hradec Kralove, Czech Republic, (with J. Plavka)

• 11:50 On the robustness of matrices in max-min algebra, Plavka, Jan, Technical University Kosice, Slovakia

• 12:15 Nonnegative linear algebra and max linear algebra: where’s the difference?, Schneider, Hans, University of Wisconsin, Madison, USA

Contributed Talks, Wednesday Morning, Room 4

• 10:35 Convex and Non-convex Optimization Problems for the Field of Values of a Matrix Uhlig, Frank, Mathematics, Auburn University, Auburn, AL 36849, USA

• 11:00 Total least-squares regularization of Tykhonov type and an ancient racetrack in Corinth Schaffrin, Burkhard, Ohio State University, Columbus, OH, USA (with Kyle Snow)

• 11:25 Numerical ranges of nilpotent operators Wu, Pei Yuan, National Chiao Tung University, Hsinchu, Taiwan (with Hwa-Long Gau)

• 11:50 p-radii of linear operators and equations of self-similarity Protasov, Vladimir, Moscow State University, Moscow, Russia

• 12:15 Matrix Optimization in Statistics Seddighin, Morteza, Indiana University East, Richmod, Indiana

Parallel Session 6, Thursday Morning

Time  MS1 Room 1  MS5 Room 2  CT Room 3  CT Room 4  CT Room 5
10:35  Barrett, Wayne  Gugercin, Serkan  Wrobel, Iwona  Sivic, Klemen  Körtesi, Peter
11:00 van der Holst, Hein Van Dooren, Paul Sánchez, Salvador Marovt, Janko  Im, Bokhee
11:25  Grout, Jason  Beattie, Christopher Furtado, Susana Olshevsky, Vadim  Al Zhour, Zeyad
11:50  Shader, Bryan  Feng, Lihong  Sendov, Hristo Palma, Alejandro  Sebeldin, Anatoly
12:15  Deaett, Louis  Freund, Roland  Fonseca, Carlos  Satô, Kenzi  Shahryari, Mohammad

MS1 Combinatorial Matrix Theory, Thursday Morning, Room 1

• 10:35 Minimum rank of edge subdivisions of complete graphs and wheels, Barrett, Wayne, Brigham Young University, Utah, (with Ryan Bowcutt, Mark Cutler, Seth Gibelyou, and Kayla Owens)

• 11:00 Computing the minimum rank of partial 2-trees, van der Holst, Hein, Eindhoven University of Technology, The Netherlands

• 11:25 Characterizing graphs with minimum rank at most a given number over a finite field using polarities of projective geometries, Grout, Jason, Iowa State University, Ames, IA, USA

• 11:50 Average minimum rank of a graph, Shader, Bryan, University of Wyoming, USA, (with Francesco Barioli, Shaun Fallat, Tracy Hall, Daniel Hershkowitz, Leslie Hogben, Ryan Martin, and Hein van der Holst)

• 12:15 The graph and rank of a positive semidefinite matrix, Deaett, Louis, University of Wisconsin, Madison, USA

MS5 Linear Algebra in Model Reduction, Thursday Morning, Room 2

• 10:35 A Krylov-Based Descent Approach for the Optimal H2 Model Reduction of Large-Scale Dynamical Systems, Gugercin, Serkan, Virginia Tech, Blacksburg, VA, USA

• 11:00 H2 Approximation and Tangential Rational Interpolation, Van Dooren, Paul, Université Catholique Louvain-La-Neuve, Belgium

• 11:25 Interpolatory Projection Methods for Parameterized Model Reduction, Beattie, Christopher, Virginia Tech, Blacksburg, VA, USA • 11:50 Parametric Model Reduction for Systems with Coupled Parameters,

Feng, Lihong, Chemnitz University of Technology, Germany/Fudan University Shanghai, China

• 12:15 The Effects of Deflation in Projection-Based Order Reduction, Freund, Roland, University of California, Davis, USA

Contributed Talks, Thursday Morning, Room 3

• 10:35 The Gauss-Lucas theorem and the numerical range Wrobel, Iwona, Warsaw University of Technology and Polish Academy of Sciences, Warsaw, Poland

• 11:00 Manifold of proper elements Sánchez Perales, Salvador, Benemérita Universidad Autónoma de Puebla, Puebla, México (with S. Djordjevic)

• 11:25 Order Invariant Spectral Properties for Several Matrices Furtado, Susana, Centro de Estruturas Lineares e Combinatórias da U. L, Porto, Portugal (with Charles Johnson)

• 11:50 Spectral Manifolds Sendov, Hristo, The University of Western Ontario, London, Canada (with A. Daniilidis, J. Malick, and A. Lewis)

• 12:15 An inequality for the multiplicity of an eigenvalue Fonseca, Carlos, Department of Mathematics, University of Coimbra, Coimbra, Portugal

Contributed Talks, Thursday Morning, Room 4

• 10:35 On varieties of commuting triples Sivic, Klemen, Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia

• 11:00 Homomorphisms of matrix semigroups over division rings from dimension two to three Marovt, Janko, Institute of Mathematics, Physics and Mechanics, Ljubljana, Slovenia (with Gregor Dolinar)

• 11:25 Can One Invert a Matrix via Graph Manipulations?

Olshevsky, Vadim, University of Connecticut, Storrs, USA (with Tom Bella and Pavel Zhlobich)

• 11:50 Solution of the linear time-dependent potential by using a solvable Lie algebra^{∗} Palma, Alejandro, Instituto de Física (BUAP), Puebla, México (with M. Villa, and L. Sandoval)

• 12:15 The algebraic relations of curvatures of PL manifolds Satô, Kenzi, Tamagawa University, Tokyo, Japan

Contributed Talks, Thursday Morning, Room 5

• 10:35 Using Linear Algebra in Teaching Hamilton Quaternions and Graphs Kortesi, Peter, University of Miskolc, Miskolc, Hungary

• 11:00 Representations of trilinear products in Comtrans algebras Im, Bokhee, Chonnam National University, Gwangju, Korea(Rep. of) (with Jonathan D. H. Smith)

• 11:25 Matrix Results on Weighted Drazin Inverse and Some Applications Al Zhour, Zeyad, Zarqa Private University, Zarqa, Jordan (with Adem Kilicman)

• 11:50 Algorithm resolving problem of determination of finite cyclicgroup by its automorphism group Sebeldin, Anatoly, University UGANC, Guinea, Conakry, Guinea (with V. K. Vildanov and A. L. Sylla)

• 12:15 Z_{2}-graded symmetry classes of tensors Shahryari, Mohammad, Tabriz University, Tabriz, Iran, Tabriz, Iran

Parallel Session 7, Thursday Afternoon

Time MS2 Room 1  MS3 Room 2  CT Room 3  CT Room 4  CT Room 5
16:55 Dhillon, Inderjit  Rust, Bert W.  Day, Jane  Marques, Maria-da-Graça Prokip, Volodymyr
17:20  Cox, Steve  Bardsley, John  Mikkelson, Rana  Dolinar, Gregor  Rakotondrajao, Fanja
17:45 Meerbergen, Karl  Nagy, James  Perdigão, Cecília  Ahn, Eunkyung
18:10 Plestenjak, Bor  Hnetynková, Iveta  Mitchell, Lon  da Cruz, Henrique F
18:35  Renaut, Rosemary A Sinkovic, John  Lee, Gue Myung

MS2 Eigenproblems: Theory and computation, Thursday Afternoon, Room 1

• 16:55 On some modified root-finding problems , Dhillon, Inderjit, University of Texas, Austin, TX, USA, (with Matyas Sustik)

• 17:20 Eigen-reduction of Large Scale Neuronal Networks , Cox, Steve, Rice University, Houston TX, USA, (with Tony Kellems, Derrick Roos and Nan Xiao)

• 17:45 Recycling Ritz vectors in the parameterized Lanczos method, Meerbergen, Karl, Katholieke Universiteit Leuven, Belgium, (with Zhaojun Bai)

• 18:10 Numerical methods for two-parameter eigenvalue problems, Plestenjak, Bor, University of Ljubljana, Slovenia

MS3 Implementation and Application issues in regularizing least squares, Thursday Afternoon, Room 2

• 16:55 A Truncated Singular Component Method for Ill-Posed Problems, Rust, Bert W. , NIST, USA, (with Dianne P. O’Leary, University of Maryland)

• 17:20 Truncation Rules for Iterative Deblurring Methods, Bardsley, John , University of Montana, USA

• 17:45 Lanczos Hybrid Regularization for Image Processing Applications, Nagy, James , Emory University, Atlanta GA, USA

• 18:10 On solvability of total least squares problem, Hnetynková, Iveta , Charles University in Prague, and Institute of Computer Science, Academy of Sciences, Czech Republic.

• 18:35 A Newton Iteration for estimating the regularizing parameter for least squares, Renaut, Rosemary A. , Arizona State University, Phoenix, USA, (with Jodi Mead, Boise State University)

Contributed Talks, Thursday Afternoon, Room 3

• 16:55 Graph Energy Change Due to Edge Deletion Day, Jane, San Jose State University, San Jose, CA, USA

• 17:20 Minimum Rank of Graphs with Loops Mikkelson, Rana, Iowa State University, Ames, IA, United States of America

• 17:45 On the equivalence class graph Perdigão, Cecília, Faculdade de Ciências e Tecnologia-UNL, Lisboa, Portugal (with Rosário Fernandes)

• 18:10 Orthogonal Removal of Vertices and Minimum Semidefinite Rank Mitchell, Lon, Virginia Commonwealth University, Richmond, United States (with Sivaram Narayan)

• 18:35 An upper bound for the maximum nullity of a symmetric matrix whose graph is outerplanar Sinkovic, John, Technische Universiteit Eindhoven, Eindhoven, Netherlands

Contributed Talks, Thursday Afternoon, Room 4

• 16:55 Some conditions for the commutativity of matrix patterns Marques, Maria-da-Graça, University of Algarve and CELC, Faro, Portugal (with C. R. Johnson)

• 17:20 General preservers of quasi-commutativity Dolinar, Gregor, Faculty of Electrical Engineering, Ljubljana, Slovenia (with Bojan Kuzma)

• 17:45 An extended Lie-Trotter formula and its applications Ahn, Eunkyung, Kyungpook National University, Daegu, Korea (with Sejong Kim and Yongdo Lim)

• 18:10 On the matrices that preserve the value of the immanant of the upper triangular matrices da Cruz, Henrique F., U.B.I, Covilhã, Portugal (with Rosário Fernandes)

• 18:35 Complexity Analysis of the Primal-Dual Interior Point Method for Second-order Cone Optimization Problem Lee, Gue Myung, Pukyong National University, Busan, Korea (with Bo Kyung Choi)

Contributed Talks, Thursday Afternoon, Room 5

• 16:55 On the problem of diagonalizability of matrices over a principal ideal domain Prokip, Volodymyr, Institute of Appl. Problem for Mech. and Math. NAS of Ukr, L’viv, Ukraine

• 17:20 Euler’s difference table and maximum permanents of (0, 1)-matrices Rakotondrajao, Fanja, University of Antananarivo, Antananarivo, Madagascar

Parallel Session 8, Friday Morning

Time MS2 Room 1  MS3 Room 2  CT Room 3  CT Room 4
10:35 Olshevsky, Vadim Estatico, Claudio  Andjelic, Milica  Dopazo, Esther
11:00  Bella, Tom  Aricó, Antonio  Martins, Enide  Bengochea, Gabriel
11:25 Li, Chi-Kwong Stefan, Wolfgang  Schaeffer, Elisa  Castro-González, Nieves
11:50 Karow, Michael  Mead, Jodi  Rodríguez, Juan Alberto  Patricio, Pedro
12:15  Moro, Julio  Narayan, Sivaram  Verde-Star, Luis

MS2 Eigenproblems: Theory and computation, Friday Morning, Room 1

• 10:35 Lipschitz stability of canonical Jordan bases of H-selfadjoint matrices, Olshevsky, Vadim, University of Connecticut, Storrs, CT, USA, (with Tom Bella and Upendra Prasad)

• 11:00 Eigenproblems for quasiseparable matrices, Bella, Tom, University of Connecticut, Storrs, CT, USA, (with Yuli Eidelman, Israel Gohberg and Vadim Olshevsky)

• 11:25 Eigenvalues of the sum of matrices from unitary similarity orbits, Li, Chi-Kwong, College of William and Mary, Williamsburg, VA, USA, (with Yiu-Tung poon and Nung-Sing Sze)

• 11:50 Pseudospectra and Stability radii for Hamiltonian Matrices, Karow, Michael, Technische Universitat Berlin, Germany

• 12:15 Structured Holder condition numbers for eigenvalues under fully nongeneric perturbations, Moro, Julio, Universidad Carlos III de Madrid, Spain, (with María J. Peláez)

MS3 Implementation and Application issues in regularizing least squares, Friday Morning, Room 2

• 10:35 Block splitting least square regularization for structured matrices arising in nonlinear microwave imaging, Estatico, Claudio , University of Cagliari, Italy

• 11:00 Signal and Image regularization via antireflective transform, Aricó, Antonio , Università degli studi di Cagliari, Italy

• 11:25 Regularizing Least Squares with the Concentration Method, Stefan, Wolfgang , Arizona State University, (with Rosemary Renaut and Anne Gelb)

• 11:50 Calculating Weights in Least Squares Estimation Using the Chi-squared Method, Mead, Jodi , Boise State University, Idaho, (with Rosemary Renaut, ASU)

Contributed Talks, Friday Morning, Room 3

• 10:35 One upper bound for the largest eigenvalue of the signless Laplacian Andjelic, Milica, Center for Research on Optimization and Control, Aveiro, Portugal (with Slobodan Simic)

• 11:00 On the spectra of some graphs like weighted rooted trees Martins, Enide, Centre for Research on Optimization and Control (CEOC), Aveiro, Portugal (with Rosário Fernandes and Helena Gomes)

• 11:25 Locally computable approximations of absorption times for graph clustering Schaeffer, Elisa, Universidad Autónoma de Nuevo León, San Nicolás de los Garza, México (with Pekka Orponen and Vanesa Avalos)

• 11:50 The Laplacian Spectrum of Hypergraphs Rodríguez, Juan Alberto, Universitat Rovira i Virgili, Tarragona, España (with Aida Kamisalic)

• 12:15 Linearly Independent Vertices and Minimum Semidefinite Rank Narayan, Sivaram, Central Michigan University, Mount Pleasant, Michigan 48859, USA

Contributed Talks, Friday Morning, Room 4

• 10:35 Further results on the representation of the Drazin inverse of a 2 × 2 block matrix Dopazo, Esther, Facultad de Informática. Universidad Politécnica de Madrid, Boadilla del Monte, Madrid, Spain (with M. F. Martínez-Serrano and N. Castro-González)

• 11:00 Duality in the Hopf Algebra of multivariate polynomials Bengochea, Gabriel, Universidad Autónoma de la Ciudad de México, México D.F. (with L. Verde-Star)

• 11:25 Representations for the generalized Drazin inverse of additive perturbations Castro-González, Nieves, Universidad Politécnica de Madrid, Madrid, Spain (with M. F. Martínez-Serrano)

• 11:50 Some additive results on Drazin Inverses Patricio, Pedro, Departamento de Matemática, Universidade do Minho, Braga, Portugal (with R. E. Hartwig)

• 12:15 Linear algebraic approach to rational functions Verde-Star, Luis, UAM Iztapalapa, México DF, México

Parallel Session 9, Friday Afternoon

Time  MS1 Room 1  MS5 Room 2  CT Room 3  CT Room 4
15:30  Olesky, Dale  Baur, Ulrike  Cortés, Vanesa  Guzmán, José Ramón
15:55 Vander Meulen, Kevin Damm, Tobias Teixeira Matos, Isabel  Milligan, Thomas
16:20  DeAlba, Luz  Benner, Peter  Goldberger, Assaf  Barría, José
16:45  Tam, Bit-Shun  Wojciechowski, Piotr  Cheng, Wei
17:10  Esen, Özlem

MS1 Combinatorial Matrix Theory, Friday Afternoon, Room 1

• 15:30 Group Inverses of Matrices with Path Graphs, Olesky, Dale, University of Victoria, Canada, (with M. Catral and P. van den Driessche)

• 15:55 Sparse Inertially Arbitrary Sign Patterns, Vander Meulen, Kevin, Redeemer College, Canada, (with L. Vanderspek and M. Cavers)

• 16:20 The Q-matrix completion problem , DeAlba, Luz, Drake University, Des Moines, IA, USA, (with Leslie Hogben and Bhaba Sharma)

• 16:45 Maximizing spectral radius of unoriented Laplacian matrix, Tam, Bit-Shun, Tamkang University, Tamsui, Taiwan, (with Ding-Jung Chang and Shui-Hei Wu)

MS5 Linear Algebra in Model Reduction, Friday Afternoon, Room 2

• 15:30 Model Reduction for unstable Systems based on Hierarchical Matrix Arithmetic, Baur, Ulrike, Chemnitz University of Technology, Germany

• 15:55 Algebraic Gramians and Model Reduction for Different System Classes, Damm, Tobias, TU Kaiserslautern, Germany

• 16:20 Balancing-Related Model Reduction for Large-Scale Unstable Systems, Benner, Peter, Chemnitz University of Technology, Germany

Contributed Talks, Friday Afternoon, Room 3

• 15:30 Some properties of the class sign regular matrices and its subclasses Cortés, Vanesa, Universidad de Zaragoza, Zaragoza, Spain (with J. M. Peña)

• 15:55 A Completion Problem over the Field of Real Numbers Teixeira Matos, Isabel, Centro de Estruturas Lineares e Combinatórias (CELC), Lisboa, Portugal (with Fernando C. Silva)

• 16:20 An upper bound on the characteristic polynomial of a nonnegative matrix leading to a proof of the Boyle– Handleman conjecture Goldberger, Assaf, Tel Aviv University, Tel Aviv, Israel (with Michael Neumann)

• 16:45 Orderings of matrix algebras and their applications Wojciechowski, Piotr, University of Texas at El Paso, El Paso, USA

• 17:10 On The Root Clustering of Matrices Esen, Özlem, Anadolu University, Eskisehir, Turkey

Contributed Talks, Friday Afternoon, Room 4

• 15:30 Reduction of an Ito’s diffusion input output model for the determination of square mean stability Guzmán, José Ramón, Instituto de Investigaciones Económicas. UNAM., México, México

• 15:55 On Euclidean Squared Distance Matrices Milligan, Thomas, University of Central Oklahoma, Edmond, OK, USA (with Chi-Kwong Li and Michael Trossett)

• 16:20 The strong closure of the similarity orbit for a class of pairs of finite rank operators Barría, José, Santa Clara University, Santa Clara, U.S.A.

• 17:10 One Type of Inverse Eigenvalue Problems in Quaternionic Quantum Mechanics Cheng, Wei, National University of Defense Technology, Changsha, P.R. China (with Liang-gui Feng)

Index

Ahn, Eunkyung: Thu. 17:45, Room 4 Al Zhour, Zeyad: Thu. 11:25, Room 5 Andjelic, Milica: Fri. 10:35, Room 3 Arav, Marina: Tue. 10:35, Room 3 Aricó, Antonio: MS3, Fri. 11:00, Room 2 Bardsley, John: MS3, Thu. 17:20, Room 2 Barrett, Wayne: MS1, Thu. 10:35, Room 1 Barría, José: Fri. 16:20, Room 4 Baur, Ulrike: MS5, Fri. 15:30, Room 2 Beattie, Christopher: MS5, Thu. 11:25, Room 2 Bella, Tom: MS2, Fri. 11:00, Room 1 Bengochea, Gabriel: Fri. 11:00, Room 4 Benner, Peter: MS5, Fri. 16:20, Room 2 Berman, Avi: MS1, Wed. 12:15, Room 1 Bini, Dario: MS6, Tue. 10:35, Room 1 Boettcher, Albrecht: Plenary, Mo. 15:30–16:25 Boimond, Jean-Louis: MS7, Tue. 16:55, Room 3 Bourgeois, Gerald: Mon. 17:20, Room 4 Brualdi, Richard A.: MS1, Wed. 10:35, Room 1 Bru, Rafael: Mon. 18:10, Room 3 Bueno Cachadina, María Isabel: Tue. 17:20, Room 4 Butkovic, Peter: MS7, Tue. 18:10, Room 3 Carriegos, Miguel: Tue. 11:50, Room 4 Castro-González, Nieves: Fri. 11:25, Room 4 Catral, Minerva: MS8, Mon. 18:35, Room 1 Cheng, Wei: Fri. 17:10, Room 4 Corral, Cristina: Mon. 11:10, Room 3 Cortés, Vanesa: Fri. 15:30, Room 3 Costa, Liliana: Mon. 10:45, Room 4 Cox, Steve: MS2, Thu. 17:20, Room 1 Cravo, Glória: Tue. 10:35, Room 4 da Cruz, Henrique F.: Thu. 18:10, Room 4 Damm, Tobias: MS5, Fri. 15:55, Room 2 Day, Jane: Thu. 16:55, Room 3 Deaett, Louis: MS1, Thu. 12:15, Room 1 DeAlba, Luz: MS1, Fri. 16:20, Room 1 Dhillon, Inderjit: MS2, Thu. 16:55, Room 1 Dodig, Marija: Mon. 10:45, Room 3 Dogan-Dunlap, Hamide: MS4, Mon. 11:10, Room 1 Dolinar, Gregor: Thu. 17:20, Room 4 Domínguez, María Elena: Mon. 17:20, Room 3 Dopazo, Esther: Fri. 10:35, Room 4 Esen, Özlem: Fri. 17:10, Room 3 Estatico, Claudio: Fri. 10:35, Room2 Eubanks, Sherod: MS8, Mon. 18:10, Room 1 Fassbender, Heike: Plenary, Thu. 8:10–9:05; MS6, Tue. 11:00, Room1 Feng, Lihong: MS5, Thu. 11:50, Room 2 Fernandes, Rosário: Mon. 11:10, Room 4 Ferrer, Josep: Tue. 12:15, Room 4 Fonseca, Carlos: Thu. 12:15, Room 3 Fošner, Ajda: Mon. 11:35, Room 3 Frank, Martin: Tue. 17:45, Room 4 Freund, Roland: MS5, Thu. 12:15, Room 2 Furtado, Susana: Thu. 11:25, Room 3 Furuichi, Shigeru: Mon. 16:55, Room 4 Gassó, Maria T.: Tue. 12:15, Room 3 Gaubert, Stephane: MS7, Mon. 10:45, Room 2 Gavalec, Martin: MS7, Wed. 11:25, Room 3 Gemignani, Luca: Plenary, Thu. 15:30–16:25 Glebsky, Lev: Tue. 11:50, Room 3 Goldberger, Assaf: Fri. 16:20, Room 3 Gouveia, María: Mon. 16:55, Room 3 Grout, Jason: MS1, Thu. 11:25, Room 1 Grudsky, Sergey: Mon. 18:35, Room 3 Gugercin, Serkan: MS5, Thu. 10:35, Room 2 Guo, Chun-Hua: MS6, Tue. 11:25, Room 1 Guzmán, José Ramón: Fri. 15:30, Room 4 Harel, Guershon: MS4, Mon. 10:45, Room 1 Hershkowitz, Danny: MS1, Wed. 11:00, Room 1 Hnetynková, Iveta: MS3, Thu. 18:10, Room 2 Hogben, Leslie: Plenary, Mo. 8:20–9:15 Horn, Roger: Mon. 12:00, Room 3 Iannazzo, Bruno: MS6, Tue. 11:50, Room 1 Im, Bokhee: Thu. 11:00, Room 5 Jiang, ErXiong: Plenary, Mo. 9:20–10:15 Karow, Michael: MS2, Fri. 11:50, Room 1 Kirkland, Steve: MS1, Wed. 11:25, Room 1 Klasa-Bompoint, Jacqueline: MS4, Tue. 16:55, Room 1 Klein, Andre: MS6, Tue. 12:15, Room 1 Kopparty, Bhaskara Rao: Mon. 18:10, Room 2 Körtesi, Peter: Thu. 10:35, Room 5 Kressner, Daniel: Plenary, Tu. 9:10–10:05 Laffey, Thomas: MS8, Mon. 17:45, Room 1 Lancaster, Peter: MS6, Tue. 16:55, Room 2 Lee, Gue Myung: Thu. 18:35, Room 4 Lee, Hosoo: Mon. 18:35, Room 4 Leon, Steven: MS4, Tue. 17:20, Room 1 Li, Chi-Kwong: MS2, Fri. 11:25, Room 1 Loiseau, Jean Jacques: MS7, Wed. 11:00, Room 3 Machado, Silvia: Mon. 16:55, Room 2 Maracci, Mirko: MS4, Mon. 11:35, Room 1 Marovt, Janko: Thu. 11:00, Room 4 Marques, Maria-da-Graça: Thu. 16:55, Room 4 Martínez, José-Javier: Tue. 18:35, Room 4 Martin, William: MS4, Tue. 18:10, Room 1 Martins, Enide: Fri. 11:00, Room 3 McDonald, Judith: MS8, Mon. 16:55, Room 1 McEneaney, William: MS7, Mon. 11:35, Room 2 M. Dopico, Froilán: Plenary, Tu. 15:30–16:25 Mead, Jodi: MS3, hd , Room 2 Meerbergen, Karl: MS2, Thu. 17:45, Room 1 Meini, Beatrice: MS6, Tue. 17:20, Room 2 Mena, Hermann: MS6, Tue. 17:45, Room 2 Merlet, Glenn: MS7, Tue. 17:45, Room 3 Mikkelson, Rana: Thu. 17:20, Room 3 Milligan, Thomas: Fri. 15:55, Room 4 Mitchell, Lon: Thu. 18:10, Room 3 Moro, Julio: MS2, Fri. 12:15, Room 1 Morris, DeAnne: MS8, Tue. 11:00, Room 2 Moura, Ana: MS4, Wed. 10:35, Room 2 Nagy, James: Plenary, Wed. 9:10–10:05; MS3, Thu. 17:45, Room 2 Narayan, Sivaram: Fri. 12:15, Room 3 Neumann, Michael: MS8, Tue. 11:50, Room 2 Olesky, Dale: MS1, Fri. 15:30, Room 1 Olshevsky, Vadim: Thu. 11:25, Room 4; MS2, Fri. 10:35, Room 1 Palma, Alejandro: Thu. 11:50, Room 4 Parraguez, Marcela: MS4, Mon. 12:00, Room 1 Patricio, Pedro: Fri. 11:50, Room 4 Peña, Juan Manuel: Plenary, Wed. 8:10–9:05 Peña, Marta: Tue. 11:00, Room 4 Perdigão, Cecília: Thu. 17:45, Room 3 Plavka, Jan: MS7, Wed. 11:50, Room 3 Plestenjak, Bor: MS2, Thu. 18:10, Room 1 Ponce, Daniela: MS7, Mon. 12:25, Room 2 Poole, George: Mon. 17:20, Room 2 Possani, Edgar: MS4, Wed. 11:25, Room 2 Prokip, Volodymyr: Thu. 16:55, Room 5 Protasov, Vladimir: Wed. 11:50, Room 4 Pruneda, Rosa E.: Tue. 11:00, Room 3 Pryporova, Olga: MS1, Wed. 11:50, Room 1 Renaut, Rosemary A.: MS3, Thu. 18:35, Room 2 Roca, Alicia: Tue. 11:25, Room 4 Rodríguez, Juan Alberto: Fri. 11:50, Room 3 Rosenthal, Peter: Plenary, Fri. 8:10–9:05 Rump, Siegfried M.: Mon. 17:45, Room 3 Russo, Maria Rosaria: Tue. 16:55, Room 4 Rust, Bert W.: MS3, Thu. 16:55, Room 2 Salam, Ahmed: Tue. 18:10, Room 4 Sánchez Perales, Salvador: Thu. 11:00, Room 3 Satô, Kenzi: Thu. 12:15, Room 4 Schaeffer, Elisa: Fri. 11:25, Room 3 Schaffrin, Burkhard: Wed. 11:00, Room 4 Schneider, Hans: MS7, Wed. 12:15, Room 3 Sebeldin, Anatoly: Thu. 11:50, Room 5 Seddighin, Morteza: Wed. 12:15, Room 4 Semrl, Peter: Tue. 11:25, Room 3 Sendov, Hristo: Thu. 11:50, Room 3 Sergeev, Sergey: MS7, Tue. 17:20, Room 3 Shader, Bryan: MS1, Thu. 11:50, Room 1 Shahryari, Mohammad: Thu. 12:15, Room 5 Shaked-Monderer, Naomi: Plenary, Tu. 8:10–9:05 Singer, Ivan: MS7, Wed. 10:35, Room 3 Sinkovic, John: Thu. 18:35, Room 3 Sivic, Klemen: Thu. 10:35, Room 4 Šmigoc, Helena: MS8, Mon. 17:20, Room 1 Soares, Graça: Mon. 17:45, Room 4 Spitkovsky, Ilya: Plenary, Fri. 9:10–10:05 Stefan, Wolfgang: MS3, Fri. 11:25, Room 2 Stosic, Marko: Mon. 12:25, Room 3 Strong, David: MS4, Wed. 11:50, Room 2 Stuart, Jeffrey: Mon. 11:35, Room 4 Szyld, Daniel: MS8, Tue. 10:35, Tue. 11:25, Room 2 Tam, Bit-Shun: MS1, Fri. 16:45, Room 1 Tanguay, Denis: MS4, Tue. 17:45, Room 1 Teixeira Matos, Isabel: Fri. 15:55, Room 3 Trigueros, María: MS4, Mon. 12:25, Room 1 Uchiyama, Mitsuru: Mon. 18:10, Room 4 Uhlig, Frank: Wed. 10:35, Room 4 Vallejo, Ernesto: Mon. 12:25, Room 4 van den Driessche, Pauline: MS7, Mon. 11:10, Room 2 van der Holst, Hein: MS1, Thu. 11:00, Room 1 Vander Meulen, Kevin: MS1, Fri. 15:55, Room 1 Van Dooren, Paul: Plenary, Thu. 9:10–10:05; MS5, Thu. 11:00, Room 2 Vargas, Xaab Nop: Mon. 17:45, Room 2 Verde-Star, Luis: Fri. 12:15, Room 4 Vieira, Luis: Mon. 18:35, Room 2 Weaver, James: Mon. 12:00, Room 4 Wojciechowski, Piotr: Fri. 16:45, Room 3 Wrobel, Iwona: Thu. 10:35, Room 3 Wu, Pei Yuan: Wed. 11:25, Room 4 Zandieh, Michelle: MS4, Wed. 11:00, Room 2 Zimmermann, Karel: MS7, Mon. 12:00, Room 2 Zúñiga, Juan Carlos: MS6, Tue. 18:10, Room 2